# Out of the Box, into the Clinic? Evaluating State-of-the-Art ASR for Clinical Applications for Older Adults

### v2

@@ 6 Limitations
Though dedicated GPUs in HPC clusters have higher bandwidths and are faster, hence the default choice in experiments like ours, unified architectures (such as provided by Apple’s Silicon M series) are receiving more attention nowadays due to their benefits in terms of latency, energy consumption and system footprint, and are recognized as efficient and competitive alternatives to dedicated GPUs (Hübner et al., 2025; Kenyon and Capano, 2022). So faster processing times could probably be attained by using dedicated GPUs or by using a more recent M-chip. Still, we anticipate a scenario where one device hosts multiple smaller AI models to do different tasks, for which our setup can provide a good lower bound.

### v3

@@ 6 Limitations
Though dedicated GPUs in high performing computing clusters have higher bandwidths and are faster, hence the default choice in experiments like ours, unified architectures (such as provided by Apple’s Silicon M series) are receiving more attention nowadays due to their benefits in terms of latency, energy consumption, and system footprint, and are recognized as efficient and competitive alternatives to dedicated GPUs (Hübner et al., 2025; Kenyon and Capano, 2022). So faster processing times could probably be attained by using dedicated GPUs or by using a more recent M-chip. Still, we anticipate a scenario where one device hosts multiple smaller AI models to do different tasks, for which our setup can provide a good lower bound.